what is the lateral surface area of a right hexagonal prism whose base is a regular hexagon with sides measuring 8 inches long and altitude measuring 6 inches tall
Answer by Alan3354(69443)

It's the perimeter of the hexagon times the height
P = 6*8
Area = 6*8*6 = 288 sq inches